# ASCoT: An Adaptive Self-Correction Chain-of-Thought Method for Late-Stage Fragility in LLMs

### v1

@@ Abstract
Chain-of-Thought (CoT) prompting has significantly advanced the reasoning capabilities of Large Language Models (LLMs), yet the reliability of these reasoning chains remains a critical challenge. A widely held “cascading failure” hypothesis suggests that errors are most detrimental when they occur early in the reasoning process. This paper challenges that assumption through systematic error-injection experiments, revealing a counter-intuitive phenomenon we term “Late-Stage Fragility”: errors introduced in the later stages of a CoT chain are significantly more likely to corrupt the final answer than identical errors made at the beginning. To address this specific vulnerability, we introduce the Adaptive Self-Correction Chain-of-Thought (ASCoT) method. ASCoT employs a modular pipeline in which an Adaptive Verification Manager (AVM) operates first, followed by the Multi-Perspective Self-Correction Engine (MSCE). The AVM leverages a Positional Impact Score function $I(k)$ that assigns different weights based on the position within the reasoning chains, addressing the Late-Stage Fragility issue by identifying and prioritizing high-risk, late-stage steps. Once these critical steps are identified, the MSCE applies robust, dual-path correction specifically to the failure parts. Extensive experiments on benchmarks such as GSM8K and MATH demonstrate that ASCoT achieves outstanding accuracy, outperforming strong baselines, including standard CoT. Our work underscores the importance of diagnosing specific failure modes in LLM reasoning and advocates for a shift from uniform verification strategies to adaptive, vulnerability-aware correction mechanisms.
Introduction
The advent of Chain-of-Thought (CoT) has endowed Large Language Models (LLMs) with remarkable multi-step reasoning capabilities (Nye et al. 2021; Kojima et al. 2022), enabling them to deconstruct complex problems into a sequence of manageable sub-problems (Yao et al. 2023; Shinn et al. 2023). Recent studies, such as the o1 from OpenAI  (Jaech et al. 2024) and DeepSeekR1  (Guo et al. 2025) have shown that extended CoT reasoning improves LLM performance. However, the success of CoT and its more verbose variants has introduced challenges: prohibitive computational costs from long reasoning chains (Chen et al. 2025) and, more critically, the inherent brittleness of the reasoning process itself (Zhang et al. 2024; Ke et al. 2025). A single flaw is often sufficient to invalidate an entire process of thought, raising significant concerns about the trustworthiness of these models (Chowdhury and Caragea 2025).
@@ Introduction
Figure 1: The top part (a) illustrates the standard CoT reasoning process, which often generates redundant outputs and lacks an effective self-correction mechanism when errors occur. The bottom part (b) depicts the ASCoT method, which incorporates a robust mechanism for self-correction, addressing potential reasoning flaws at each stage.
A widely held belief within the research community, often termed the “cascading failure” hypothesis, posits that the primary threat to CoT integrity lies in early-stage errors (Tyen et al. 2023). The logic is intuitive: an initial mistake in problem comprehension or a miscalculation will inevitably propagate through the logical dependency graph of the reasoning chain, contaminating all subsequent steps. Consequently, much of the research on improving CoT robustness has implicitly focused on ensuring the correctness of the initial steps through more robust prompting or post-hoc verification (Liu et al. 2024c; Yi, Wang, and Li 2025).
@@
In this work, we directly investigate this foundational assumption, seeking to answer a critical question: Are all errors in a CoT created equal in their impact? To find out this, we conduct a systematic error-injection experiment, which introduces controlled symbolic and numeric errors at varying positions within correct reasoning chains. We then instruct the LLM to resume the reasoning process from the point at which the error occurred. Our findings are striking and counter-intuitive. We discover that errors introduced in later stages are more likely to lead to an incorrect final answer. We term this phenomenon “Late-Stage Fragility”. An early error often appears to trigger the model’s latent self-correction mechanisms, allowing it to recover and still reach the correct solution. However, as the model progresses steadily, its reasoning process gains a form of “semantic commitment”, reducing relevant information entropy and flexibility. At the later stages, the model is unable to evaluate the validity of the final calculations.
@@
This empirical discovery suggests that a significant portion of efforts to secure CoT reasoning may be misdirected. As illustrated in Figure  1(a), a standard CoT process can be both inefficient due to redundancy and unreliable due to its inability to correct these late-stage errors. Building on a rethinking of how we evaluate and how we correct LLM reasoning, we introduce the ASCoT method. In contrast to standard methods, ASCoT is designed to address both challenges simultaneously. As shown in Figure  1(b), it first employs a semantic pruning stage to enhance efficiency, followed by a targeted, risk-aware correction mechanism that focuses computational resources precisely on the most vulnerable steps to ensure a correct final answer.
We evaluate ASCoT across various models, including the LLaMA-3.1-8B-Instruct and the Qwen2.5-Instruct series (3B, 7B, and 14B), using two widely recognized reasoning benchmarks: GSM8K and MATH.
@@
Efficient CoT in LLM Reasoning
A critical question for CoT efficiency is whether every token in different position contributes equally to the final answer (Xia et al. 2025). Several early approaches (Hou et al. 2022; Zhang et al. 2023; Chen et al. 2024a) measure importance based on the self-information or surprisal from a unidirectional language model:
@@
To address this, more advanced methods like LLMLingua-2 (Pan et al. 2024) utilize a bidirectional language model (Jawahar, Sagot, and Seddah 2019; Devlin et al. 2019) (e.g., BERT-style) trained specifically for token classification. This approach measures importance by the predicted probability of a token being essential (Achiam et al. 2023), conditioned on the entire sequence:
$I_{2}(x_{i})=P(x_{i}|x_{<n};\theta_{\mathcal{M}_{L}}),\quad$ (2)
@@
where $\theta_{\mathrm{\mathcal{M}_{L}}}$ represents the parameters of the trained bidirectional model. Such methods confirm that tokens within a CoT sequence are highly variable in importance. For example, tokens comprising mathematical equations or key entities (Ma et al. 2025) are identified as critical, whereas semantic connectors such as “so” and “since” contribute much less to the final outcome of the reasoning  (Hou et al. 2025). This highlights a significant potential for enhancing CoT efficiency by pruning these redundant tokens.
Error Propagation in Reasoning Process
@@
Beyond efficiency, the primary challenge for CoT is reliability. A single error can invalidate an entire reasoning process (Guo et al. 2017; Tyen et al. 2023). The conventional understanding of this problem is rooted in the “cascading failure” hypothesis, which posits that errors are most damaging when they occur early in the reasoning chain (Jiang et al. 2025; Neumann and Singh 2025). According to this view, an early mistake propagates through the logical dependency graph, contaminating all subsequent steps and ensuring an incorrect outcome. In this study, our preliminary investigations reveal a counter-intuitive dynamic that challenges this assumption. Through systematic error-injection experiments, we observe that errors introduced in the later stages of a reasoning chain are significantly more detrimental to the final outcome than identical errors introduced at the beginning. We term this phenomenon “Late-Stage Fragility”.
Method
@@
Figure 2: Illustration of ASCoT Pipeline. ASCoT generates the $CoT_{initial}$ from the target LLM, which is then compressed by the IRM module to a predefined ratio $\gamma$. The compressed CoT is verified by the AVM module, and if the confidence score exceeds a threshold $\tau$, the problematic steps are sent to the MSCE module for error correction. The corrected $CoT_{final}$ is then used for fine-tuning. Finally, ASCoT enables reasoning on new problems with the specified compression ratio $\gamma$.
@@
In this section, we detail the architecture of our method. ASCoT generates the $CoT_{initial}$ based on the target LLM. The $CoT_{initial}$ is then processed through the IRM module, reducing the output to a fixed ratio $\gamma$. The remaining CoT is passed through the AVM module, which performs a rigorous verification process. If the model’s risk score for any CoT step exceeds a predefined threshold $\tau$, indicating a potential error, the corresponding CoT step is forwarded to the MSCE module. The MSCE module applies internal and external validation techniques to correct errors, resulting in the corrected $CoT_{final}$. Ultimately, the refined, non-redundant $CoT_{final}$ is used for fine-tuning. We illustrate the training and inference process in Figure  2, respectively.
Intelligent Routing Mechanism
@@
The IRM serves as the initial processing stage, responsible for preliminary CoT pruning and compression. The core idea is inspired by methods TokenSkip (Xia et al. 2025), which recognize that not all tokens in a CoT are equally important for the final reasoning outcome. The mechanism operates by first generating a complete CoT path for a given problem. Then it employs a trained token importance model to assign a semantic importance score to each token in the chain. Based on a predefined compression ratio $\gamma$, tokens with scores below the corresponding threshold are pruned, resulting in a compressed CoT, namely the final answer.
Figure 3: The details of verification and self-correction. AVM and MSCE working together to correct a faulty step $t_{k}$. AVM computes $R(t_{k})$. MSCE then applies intrinsic and extrinsic correction to produce the final corrected $CoT_{final}$.
Adaptive Verification Manager
@@
Confidence Assessment Score
To provide a holistic measure of the single reasoning step’s quality, we first define a comprehensive quality score $Q(t_{k})$, which is a composite metric that evaluates a step $t_{k}$ across four distinct dimensions: Logical Validity $V(t_{k})$, Factual Support $G(t_{k})$, Semantic Clarity $C(t_{k})$ and Process Utility $U(t_{k})$.
@@
Factual support $G(t_{k})$ measures the correctness of all $n$ verifiable steps:
@@
where $T(a_{i})\rightarrow\{0,1\}$ is the verification function that evaluates whether the $i$-th action $a_{i}$ correctly invokes the appropriate external tool.
Semantic clarity $C(t_{k})$ quantifies the clarity of step $t_{k}$ based on the average conditional probability of its tokens $\{y_{1},...,y_{m}\}$, conditioned on the prior context:
$C(t_{k})=\frac{1}{m}\sum_{i=1}^{m}\log P(y_{i}|c_{<i}),$ (5)
where $y_{i}$ is the $i$-th token of step $t_{k}$, $m$ is the total number of tokens, $c_{<i}$ is the preceding context of $y_{i}$ and $P(\cdot\mid\cdot)$ is the conditional probability produced by the LLM.
Process utility $U(t_{k})$ measures the contribution of step $t_{k}$ to the final answer $\mathcal{A}^{*}$. It is defined as the information gain, measured by the increase in the probability of the model generating $\mathcal{A}^{*}$ when $t_{k}$ is incorporated into the context:
@@
where $\mathcal{A}^{*}$ is the ground‐truth answer, $S_{\leq k}$ is the full context including step $t_{k}$ and $S_{<k}$is the context excluding step $t_{k}$.
Based on these dimensions, we design the single-step comprehensive quality scoring function $Q(t_{k})$ to integrate the multi-dimensional evaluation into a quantitative score. The function’s core design principle is a veto mechanism, implemented through a multiplicative structure. Logical validity and factual support are foundational. If a step fails on either of these criteria, it should be considered invalid regardless of its other attributes. Other assessment criteria play an auxiliary role. Therefore, the function is defined as:
$Q(t_{k})=V(t_{k})\,G(t_{k})\,\Bigl{(}w_{c}\,\frac{C(t_{k})}{C_{\max}}+w_{u}\,U(t_{k})\Bigr{)},$ (7)
where $w_{c}$ and $w_{u}$ are hyperparameters that represent the weight of semantic clarity and the weight for low process utility. Figure  3 illustrates this evaluation process, where an initial CoT with a calculation error is analyzed.
Positional Impact Score
To model “Late-Stage Fragility”, we introduce the Position Impact Function $I(k)$, which quantifies the influence of position on the system’s stability as it situates the different-stage regime:
$I(k)=w_{a}\cdot e^{\alpha(k-1)},$ (8)
where $k$ is the step’s position, $w_{a}$ is a base impact coefficient, and $\alpha$ is a propagation rate. These parameters are empirically fitted through controlled error-injection experiments. The detailed discussion of error propagation via controlled injection can be found in the experiment section. This procedure involves systematically injecting single errors at different steps $k$ in a set of baseline correct reasoning chains and measuring the probability of final answer failure. The function $I(k)$ is then fitted to these empirical data points. Finally, AVM combines these factors into a single Risk Score:
@@
Multi-Perspective Self-Correction Engine
Once a step $t_{k}$ is flagged by the AVM, the MSCE is activated to perform a robust and efficient correction. To counteract a model struggles to correct its own mistakes when using a single line of reasoning, the MSCE employs a dual-path correction strategy: Intrinsic Correction (Self-Reflection), the model is presented with the context up to step $s_{(k-1)}$ and its own flawed step $t_{k}$, and is explicitly prompted to review and amend it. Extrinsic Correction (Independent Generation), the model is given the context up to $s_{(k-1)}$ but without the flawed step $t_{k}$. It is then prompted to generate a new subsequent step from this “clean slate”.
Experiment
Experimental Setup
Models and Datasets
We primarily evaluate our method using the LLaMA-3.1-8B-Instruct (Ge et al. 2023) and Qwen2.5-Instruct series (3B, 7B, 14B) (Team 2024)models. Our evaluation leverages two notably standard math reasoning benchmarks: GSM8K (Cobbe et al. 2021), which consists of 8.5K grade school math problems requiring 2-8 reasoning steps, and MATH (Hendrycks et al. 2021), a more challenging dataset of 12.5K competition-level math problems. Given the significant computational cost associated with the MATH dataset, we follow recent research practices by evaluating our method on a carefully selected subset, MATH-500 (Lightman et al. 2023), which includes a representative sample of 500 problems.
Implementation Details
The ASCoT is implemented in PyTorch 1.13.0 and all experiments are conducted on a server with four NVIDIA GeForce RTX 3090 GPUs (24GB). For the IRM module, we use LLMLingua-2 (Pan et al. 2024) as the token importance metric to generate compressed CoT data for fine-tuning. To optimize the fine-tuning process, we adopt LoRA (Hu et al. 2022) for all fine-tuning procedures, setting the rank $r=8$ and scaling parameter $\alpha=16$ to ensure efficient training. Additional implementation details and configurations can be found in the Appendix.
@@
Table 1: Experimental results of ASCoT on LLaMA-3.1-8B-Instruct. We report Accuracy, average CoT token count (Tokens) and actual compression ratio (ActRatio) for comparison. The compression ratio $\gamma$ is set to 0.51.00.5-1.0, respectively.
Evaluation Metrics

### v3

@@ Abstract.
Chain-of-Thought (CoT) prompting has significantly advanced the reasoning capabilities of Large Language Models (LLMs), yet the reliability of these reasoning chains remains a critical challenge. A widely held “cascading failure” hypothesis suggests that errors are most detrimental when they occur early in the reasoning process. This paper challenges that assumption through systematic error-injection experiments, revealing a counter-intuitive phenomenon we term “Late-Stage Fragility”: errors introduced in the later stages of a CoT chain are significantly more likely to corrupt the final answer than identical errors made at the beginning. To address this specific vulnerability, we introduce the Adaptive Self-Correction Chain-of-Thought (ASCoT) method. ASCoT employs a modular pipeline in which an Adaptive Verification Manager (AVM) operates first, followed by the Multi-Perspective Self-Correction Engine (MSCE). The AVM leverages a Positional Impact Score function that assigns different weights based on the position within the reasoning chains, addressing the Late-Stage Fragility issue by identifying and prioritizing high-risk, late-stage steps. Once these critical steps are identified, the MSCE applies robust, dual-path correction specifically to the failure parts. Extensive experiments on benchmarks demonstrate our superiority over previous methods. ASCoT reduces token usage by 21%/30% for the LLaMA-3.1-8B on GSM8K/MATH-500 benchmarks, with minimal accuracy drops of only 1.1% and 1.8%, achieving new sota performance.
Key words and phrases:
Large Language Models, Common-Sense Reasoning, Interpretability and Evaluation of NLP Models, Prompt Engineering
1. Introduction
@@
Figure 1. The top part (a) illustrates the standard CoT reasoning process, which often generates redundant outputs and lacks an effective self-correction mechanism when errors occur. The bottom part (b) depicts the ASCoT method, which incorporates a robust mechanism for self-correction, addressing potential reasoning flaws at each stage.
The advent of Chain-of-Thought (CoT) has endowed Large Language Models (LLMs) with remarkable multi-step reasoning capabilities Kojima et al. (2022), enabling them to deconstruct complex problems into a sequence of manageable sub-problems Yao et al. (2023); Shinn et al. (2023). Recent studies, such as the o1 from OpenAI Jaech et al. (2024) and DeepSeekR1 Guo et al. (2025) have shown that extended CoT reasoning improves LLM performance. However, the success of CoT and its variants has introduced challenges: prohibitive computational costs from long reasoning chains Chen et al. (2025) and the inherent brittleness of the reasoning process itself Zhang et al. (2024); Ke et al. (2025). A single flaw is often sufficient to invalidate an entire process of thought, raising significant concerns about the trustworthiness of these models Chowdhury and Caragea (2025).
@@
A widely held belief within the research community, often termed the “cascading failure” hypothesis, posits that the primary threat to CoT integrity lies in early-stage errors Tyen et al. (2023). The logic is intuitive: an initial mistake in problem comprehension or a miscalculation will inevitably propagate through the logical dependency graph of the reasoning chain, contaminating all subsequent steps. In this work, we directly investigate this foundational assumption, seeking to answer a critical question: Are all errors in a CoT created equal in their impact? To find out this, we conduct a systematic error-injection experiment, which introduces controlled symbolic and numeric errors at varying positions within correct reasoning chains. We then instruct the LLM to resume the reasoning process from the point at which the error occurred. Our findings are striking and counter-intuitive. We discover that errors introduced in later stages are more likely to lead to an incorrect final answer. We term this phenomenon “Late-Stage Fragility”. An early error often appears to trigger the model’s latent self-correction mechanisms, allowing it to recover and still reach the correct solution.
@@
As the model progresses steadily, its reasoning process gains a form of “semantic commitment”, reducing relevant information entropy. At the later stages, the model is unable to evaluate the validity of the final calculations. This empirical discovery suggests that a significant portion of efforts to secure CoT reasoning may be misdirected. As illustrated in Figure 1(a), a standard CoT process can be both inefficient due to redundancy and unreliable due to its inability to correct these late-stage errors. Building on a rethinking of how we evaluate and how we correct LLM reasoning, we introduce the ASCoT method. In contrast to standard methods, ASCoT is designed to address both challenges simultaneously. As shown in Figure 1(b), it first employs a semantic pruning stage to enhance efficiency, followed by a targeted, risk-aware correction mechanism that focuses computational resources precisely on the most vulnerable steps to ensure a correct final answer.
@@ 2. Preliminary
Efficient CoT in LLM Reasoning A critical question for CoT efficiency is whether every token in different position contributes equally to the final answer Xia et al. (2025). Several early approaches Hou et al. (2022); Zhang et al. (2023); Chen et al. (2024b) measure importance based on the self-information or surprisal from a unidirectional language model:
@@
To address this, more advanced methods like LLMLingua-2 Pan et al. (2024) utilize a bidirectional language model Jawahar et al. (2019); Devlin et al. (2019) (e.g., BERT-style) to reframe the task as token classification. In this paradigm, the importance of a token $x_{i}$ is defined as the probability that a trained classifier assigns it an ”essential” label ($y_{i}=1$), conditioned on the entire sequence $x$:
$I_{2}(x_{i})=P(y_{i}=1|x;\theta_{\mathcal{M}{cls}}),$ (2)
@@
where $\theta_{\mathcal{M}{cls}}$ represents the parameters of the trained bidirectional classifier. Such methods confirm that tokens within a CoT sequence are variable in importance. For example, tokens comprising mathematical equations or key entities Ma et al. (2025) are identified as critical, whereas semantic connectors such as so” and since” contribute less to the final outcome of the reasoning Hou et al. (2025). This highlights that CoT efficiency can be improved by pruning these redundant tokens.
@@
Error Propagation in Reasoning Process Beyond efficiency, the primary challenge for CoT is reliability. A single error can invalidate an entire reasoning process Guo et al. (2017); Tyen et al. (2023). The conventional understanding of this problem is rooted in the “cascading failure” hypothesis, which posits that errors are most damaging when they occur early in the reasoning chain Jiang et al. (2025). According to this view, an early mistake propagates through the logical dependency graph, contaminating all subsequent steps. In this study, through systematic error-injection experiments, we observe that errors introduced in the later stages of a reasoning chain are significantly more detrimental to the final outcome than identical errors introduced at the beginning.
@@
Figure 2. Illustration of ASCoT Pipeline. ASCoT generates the $CoT_{initial}$ from the target LLM, which is then compressed by IRM module to a predefined ratio $\gamma$. The compressed CoT is verified by AVM module, and if the confidence score exceeds a threshold $\tau$, the problematic steps are sent to MSCE module for error correction. The corrected $CoT_{final}$ is then used for fine-tuning. Finally, ASCoT enables reasoning on new problems with the specified compression ratio $\gamma$.
Algorithm 1 ASCoT Mechanism
Input: Initial chain-of-thought (CoT) $CoT_{initial}$
Output: Corrected $CoT_{final}$
[…]
3:for each step $t_{k}$ in $CoT_{pruned}$ do
4:  $is\_flagged\leftarrow\mathrm{AVM}(context,t_{k})$
5:  if is_flaggedis\_flagged then
6:   $t_{k}^{\mathrm{corr}}\leftarrow\mathrm{MSCE}(context,t_{k})$
7:   append $t_{k}^{\mathrm{corr}}$ to $CoT_{final}$
8:   $context\leftarrow context\,\|\,t_{k}^{\mathrm{corr}}$
9:  else
10:   append $t_{k}$ to $CoT_{final}$
11:   $context\leftarrow context\,\|\,t_{k}$
12:  end if
13:end for
14:return $CoT_{final}$
3. Method
@@
In this section, we detail the architecture of our method. ASCoT generates the $CoT_{initial}$ based on the target LLM. The $CoT_{initial}$ is then processed through the IRM module, reducing the output to a fixed ratio $\gamma$. The remaining CoT is passed through the AVM module, which performs a rigorous verification process. If the model’s risk score for any CoT step exceeds a predefined threshold $\tau$, indicating a potential error, the corresponding CoT step is forwarded to the MSCE module. The MSCE module applies internal and external validation techniques to correct errors, resulting in the corrected $CoT_{final}$. Ultimately, the refined, non-redundant $CoT_{final}$ is used for fine-tuning. We illustrate the training and inference process in Figure 2, respectively. We also provide Pseudocode 1 for ASCoT.
3.1. Intelligent Routing Mechanism
@@
The IRM serves as the initial processing stage, responsible for preliminary CoT pruning and compression. The core idea is inspired by methods TokenSkip Xia et al. (2025), which recognize that not all tokens in a CoT are equally important for the final reasoning outcome. The mechanism operates by first generating a complete CoT path for a given problem. Then it employs a trained token importance model to assign a semantic importance score to each token in the chain. Based on a predefined compression ratio $\gamma$, tokens with scores below the corresponding threshold are pruned, resulting in a compressed CoT, namely the final answer.
3.2. Adaptive Verification Manager
@@
Confidence Assessment Score To provide a holistic measure of the single reasoning step’s quality, we first define a comprehensive quality score $Q(t_{k})$, which is a composite metric that evaluates step $t_{k}$ across four dimensions: Logical Validity $V(t_{k})$, Factual Support $G(t_{k})$, Semantic Clarity $C(t_{k})$ and Process Utility $U(t_{k})$.
@@
Factual support $G(t_{k})$ measures the correctness of all $n$ verifiable assertions within the step. For the mathematical reasoning benchmarks used in this work, verifiable assertions are arithmetic calculations. The function is defined as:
@@
where $a_{i}$ is the $i$-th arithmetic operation in step $t_{k}$, and $T(a_{i})\rightarrow\{0,1\}$ is a verification function that executes the operation using a computational engine (e.g., a Python interpreter) to confirm its accuracy.
Semantic clarity $C(t_{k})$ quantifies the coherence of step $t_{k}$. To ensure the score is well-calibrated and bounded within a $[0,1]$ range, we normalize the average conditional log-probability of its tokens $\{y_{1},...,y_{m}\}$ using a sigmoid function:
$C(t_{k})=\sigma\left(\frac{1}{m}\sum_{i=1}^{m}\log P(y_{i}|c_{<i})\right),$ (5)
where $\sigma(\cdot)$ is the sigmoid function, $y_{i}$ is the $i$-th token of step $t_{k}$, and $c_{<i}$ is the preceding context.
Process utility $U(t_{k})$ measures the contribution of step $t_{k}$ to the final answer $\mathcal{A}^{*}$. It is defined as the information gain:
@@
where $\mathcal{A}^{*}$ is the ground‐truth answer, $S_{\leq k}$ is the full context including step $t_{k}$ and $S_{<k}$is the context excluding step $t_{k}$. $U(t_{k})$ relies on the ground-truth answer and is used during the training phase. At inference time, this term is omitted.
Based on these dimensions, we design the single-step comprehensive quality scoring function $Q(t_{k})$ with a veto mechanism for foundational criteria:
$Q(t_{k})=V(t_{k})\cdot G(t_{k})\cdot\left(w_{c}\cdot C(t_{k})+w_{u}\cdot U(t_{k})\right),$ (7)
where $w_{c}$ and $w_{u}$ are hyperparameters that balance the contribution of clarity and utility. Figure 3 illustrates this evaluation process, where an initial CoT with a calculation error is analyzed.
Figure 3. The details of verification and self-correction. AVM and MSCE working together to correct a faulty step $t_{k}$. AVM computes $R(t_{k})$. MSCE then applies intrinsic and extrinsic correction to produce the final corrected $CoT_{final}$.
Positional Impact Score To quantitatively model the “Late-Stage Fragility” phenomenon observed in our error-injection experiments (Section 4.3), we introduce the Position Impact Function $I(k)$. This function is an empirical model fitted to our experimental data, designed to capture the escalating impact of errors at later reasoning stages. We normalize this position function as:
$I(k)=w_{a}\cdot e^{\alpha(k/K)},$ (8)
where $k$ is the step’s index, $K$ is the total number of steps, $w_{a}$ is a base impact coefficient, and $\alpha$ is a propagation rate. These parameters are determined by fitting the function to our error-injection data. This procedure involves injecting single errors at different steps $k$ in a set of baseline correct reasoning chains and measuring the probability of final answer failure. The function $I(k)$ is then fitted to these empirical data points. Details of the analysis are available in Appendix.
Finally, AVM combines these factors into a single Risk Score:
@@
Table 1. Experimental results of ASCoT on LLaMA-3.1-8B-Instruct. We report Accuracy, average CoT token count (Tokens) and actual compression ratio (ActRatio) for comparison. The compression ratio $\gamma$ is set to 0.51.00.5-1.0, respectively.
@@
3.3. Multi-Perspective Self-Correction Engine
Once a step $t_{k}$ is flagged by the AVM, the MSCE is activated to perform a robust and efficient correction. To counteract a model struggles to correct its own mistakes when using a single line of reasoning, the MSCE employs a dual-path correction strategy: Intrinsic Correction (Self-Reflection), the model is presented with the context up to step $s_{(k-1)}$ and its own flawed step $t_{k}$, and is explicitly prompted to review and amend it. Extrinsic Correction (Independent Generation), the model is given the context up to $s_{(k-1)}$ but without the flawed step $t_{k}$. Upon generating two candidate corrections, the MSCE employs a selection mechanism to choose the superior one. We re-apply the Confidence Assessment Score function $Q(\cdot)$ (from Eq. 7, with $U=0$) to both candidates. The candidate with the higher quality score is then integrated into the reasoning chain.
4. Experiment
4.1. Experimental Setup
Models and Datasets We primarily evaluate our method using the LLaMA-3.1-8B-Instruct Ge et al. (2023) and Qwen2.5-Instruct series (3B, 7B, 14B) Team (2024)models. Our evaluation leverages two notably standard math reasoning benchmarks: GSM8K Cobbe et al. (2021), which consists of 8.5K grade school math problems requiring 2-8 reasoning steps, and MATH Hendrycks et al. (2021), a more challenging dataset of 12.5K competition-level math problems. Given the significant computational cost associated with the MATH dataset, we follow recent research practices by evaluating our method on a carefully selected subset, MATH-500 Lightman et al. (2023), which includes a representative sample of 500 problems.
Implementation Details The ASCoT is implemented in PyTorch 1.13.0 and all experiments are conducted on a server with four NVIDIA GeForce RTX 3090 GPUs (24GB). For the IRM module, we use LLMLingua-2 Pan et al. (2024) as the token importance metric to generate compressed CoT data for fine-tuning. To optimize the fine-tuning process, we adopt LoRA Hu et al. (2022) for all fine-tuning procedures, setting the rank $r=8$ and scaling parameter $\alpha=16$ to ensure efficient training. Additional implementation details and configurations can be found in the Appendix.